<e> find the turning points and point of inflection on the curve, y=x5-5x4+5x3-1




1
Expert's answer
2022-02-25T08:50:55-0500
"y'=5x^4-20x^3+15x^2"

Find the ctitical number(s)


"y'=0=>5x^4-20x^3+15x^2=0"

"5x^2(x^2-4x+3)=0"

"5x^2 (x-1)(x-3)=0"

"x_1=0, x_2=1, x_3=3"

Find the second derivative


"y''=20x^3-60x^2+30x"

"y''(0)=0"

The function "f" has no turning point at "x=0."



"y''(1)=20-60+30=-10<0"

The function "f" has turning point at "x=1." The function "f" has a local maximum at "x=1."



"y''(3)=540-540+90=90>0"

The function "f" has turning point at "x=3." The function "f" has a local minimum at "x=3."


Find the inflection point(s)


"y''=0=>20x^3-60x^2+30x=0"

"10x(2x^2-6x+3)=0"

"5x(2x-3-\sqrt{3})(2x-3+\sqrt{3})"

Point of inflection at "x=0, x=\dfrac{3-\sqrt{3}}{2}, x=\dfrac{3+\sqrt{3}}{2}."